OpenStudy (anonymous):

Find the gravitational field strength at 220km above the surface of the earth

OpenStudy (anonymous):

You can use the equation: g=GM/r^2 to find g at any point. G=Universal Constant of Gravitation (6.67x10^-11 Nm^2/kg^2), M=mass of planet, r=radius. The mass of the Earth is 5.98x10^24 kg g = GM/r^2 g= (6.67x10^-11 Nm^2/kg^2)(5.98x10^24kg)/(6.38x10^6 m + 220,000 m)^2 [6.38x10^6m is the radius of Earth + our height above Earth, 220,000m] g = 9.16 m/s/s